# Objaverse++: Curated 3D Object Dataset with Quality Annotations

### v1

@@ 3.3.2 Classifier Validation
Among all the metrics, it is clear that the metrics for the binary tags (is_multiple_object, is_scene, is_figure, is_transparent) demonstrate strong accuracies. These metrics highlight the reliability of the network in identifying different characteristics of the model, which is valuable for various 3D modeling applications. “Score” (marked with * in Table 1) is a relatively weak metric of the annotation network due to subjective nuances, as mentioned above. Here, we include a “relaxed accuracy for score” that allows scores 2 and 3 to be interchangeable. The “relaxed score accuracy” improves significantly to 82.21%, showing that the network captures general quality distinctions effectively, though missing subtle differences among models of high quality.
@@ 4 Dataset Evaluation
To assess the performance of the model, we conducted a user study with 47 participants. Each participant was presented with 10 pairs of generation results: one generated from Training Set A and the other from Training Set B. The generated 3D models are inferenced given the sample inputs of the open-source OpenLRM. Participants were asked to choose their preferred result or to select “no preference” without knowing the origin of each generation. To reduce bias, the question order and the option order were randomized. The participants came from diverse backgrounds, including artists, machine learning researchers, game developers, and software engineers.
Figure 6: User Study Results. Of the 10 pairs of objects, 8 showed a preference for our model over the baseline. For pairs like the one titled “girl,” more than 95% of the participants chose our result. For “mailbox”, “hydrant” and “lamp”, despite the presence of the no preference option, the majority of the participants chose our generation, proving the quality is significantly higher than baseline.
@@ 4.2 Faster Convergence in Training
Quality Over Random Sampling. To ensure that dataset size is not the primary reason for faster convergence, we randomly sampled 50,000 objects from Training Set A to create a subset of comparable size to our “high-quality dataset”. As illustrated in Fig.8, the random 50k subset does not result in a significantly faster convergence than the baseline dataset. This implies that simply reducing the dataset size does not guarantee faster convergence. In comparison, our “high-quality dataset”, similar in size to the random 50k subset, produces substantially faster convergence, demonstrating that the curated quality of the samples is crucial.
Impact of Annotation Quality. The “superior quality subset”, containing only top quality samples, converges faster than the “high quality subset” dataset, which includes both high and superior quality models. This finding supports the idea that our scoring rubrics are efficient, as higher-quality data directly contributes to faster and more efficient model training.
@@ 5 Conclusion
Future directions. We acknowledge that further experiments comparing Objaverse++ with Objaverse-XL are important.
In addition, we will explore comprehensive and quantitative filtering criteria, possibly leveraging more advanced annotation networks to tag additional model attributes, such as structural complexity or aesthetic style.

### v2

@@ 3.3.2 Classifier Validation
Among all the metrics, it is clear that the metrics for the binary tags (is_multiple_object, is_scene, is_figure, is_transparent) demonstrate strong accuracies. These metrics highlight the reliability of the network in identifying different characteristics of the model, which is valuable for various 3D modeling applications. “Score" (marked with * in Table 1) is a relatively weak metric of the annotation network due to subjective nuances, as mentioned above. Here, we include a “relaxed accuracy for score" that allows scores 2 and 3 to be interchangeable. The “relaxed score accuracy" improves significantly to 82.21%, showing that the network captures general quality distinctions effectively, though missing subtle differences among models of high quality.
@@ 4 Dataset Evaluation
To assess the performance of the model, we conducted a user study with 47 participants. Each participant was presented with 10 pairs of generation results: one generated from Training Set A and the other from Training Set B. The generated 3D models are inferenced given the sample inputs of the open-source OpenLRM. Participants were asked to choose their preferred result or to select “no preference" without knowing the origin of each generation. To reduce bias, the question order and the option order were randomized. The participants came from diverse backgrounds, including artists, machine learning researchers, game developers, and software engineers.
Figure 6: User Study Results. Of the 10 pairs of objects, 8 showed a preference for our model over the baseline. For pairs like the one titled “girl," more than 95% of the participants chose our result. For “mailbox", “hydrant" and “lamp", despite the presence of the no preference option, the majority of the participants chose our generation, proving the quality is significantly higher than baseline.
@@ 4.2 Faster Convergence in Training
Quality Over Random Sampling. To ensure that dataset size is not the primary reason for faster convergence, we randomly sampled 50,000 objects from Training Set A to create a subset of comparable size to our “high-quality dataset". As illustrated in Fig.8, the random 50k subset does not result in a significantly faster convergence than the baseline dataset. This implies that simply reducing the dataset size does not guarantee faster convergence. In comparison, our “high-quality dataset", similar in size to the random 50k subset, produces substantially faster convergence, demonstrating that the curated quality of the samples is crucial.
Impact of Annotation Quality. The “superior quality subset", containing only top quality samples, converges faster than the “high quality subset" dataset, which includes both high and superior quality models. This finding supports the idea that our scoring rubrics are efficient, as higher-quality data directly contributes to faster and more efficient model training.
@@ 5 Conclusion
Future directions. We acknowledge that further experiments comparing Objaverse++ with Objaverse-XL are important. In addition, we will explore comprehensive and quantitative filtering criteria, possibly leveraging more advanced annotation networks to tag additional model attributes, such as structural complexity or aesthetic style.